Carmina is packing 27 superballs in 3 square layers. Each ball has diameter 4 cm. What is the minimum volume of the box? What is the surface area of the box? How much empty space is in the box?
Answer by richwmiller(17219)

V=(4/3)*pi*r^3
V=(4/3)*pi*2^3=32 *pi/3
The box is 3*4 =12 cms wide and long and high
V=12*12*12=144*12
Surface area is 6*12*12
Empty space is 144*12-32pi/3